OpenStudy (anonymous):

Sarah wants to deposit $2,125 into savings accounts at three different banks: Bank of the US, 7/4 Bank, and Catch Bank. She will deposit four times as much into Catch Bank as 7/4 Bank because they offer a higher interest rate. She also expects the Bank of the US deposit to be 25% of the total of the other two deposits. How much will Sheila deposit into the Bank of the US savings account? $340 $425 $515 $1,360

OpenStudy (anonymous):

I got 425 but I feel like maybe I did something wrong.

OpenStudy (anonymous):

x = Bank of US y = 7/4 bank z = Catch bank x + y + z = 2125 x = sum of the other two divided by 4. (y + 4y)/4 y = y z = 4y (4y + y) + (4y+y)/4 = 2125 5y + 5y/4 = 2125 y = 340 x = (340+340*4)/4 = 425 z = 4*340 = 1360 yes she did deposit $425 into the Bank of US.

OpenStudy (anonymous):

Okay that looks like my work too. Thank you so much!!

OpenStudy (anonymous):

you're welcome